State the domain, range, and y-intercept of points(1.6,5) (-3,1)


1
Expert's answer
2021-06-22T03:37:32-0400

Gradient , "m= \frac{\triangle y}{\triangle x}= \frac{5-1}{1.6-(-3)}= \frac{20}{23}"

Equation "\frac{20}{23}=\frac{y-1}{x+3} \implies y= \frac{20}{23}x+\frac{83}{23}"

Therefore, Domain "\left(-\infty \:,\:\infty \:\right)"

Range "\:\left(-\infty \:,\:\infty \:\right)"

y-intercept of points "\left(0,\:\frac{83}{23}\right)"